# Synthesising Activity Participations and Scheduling with Deep Generative Machine Learning

### v2

@@ Abstract
Using a deep generative machine learning approach, we synthesise human activity participations and scheduling (the choices of what activities to participate in and when). Activity schedules, which represent what people do and when, are a core component of many applied transport, energy, and epidemiology models. Our data-driven approach learns the distributions resulting from human preferences and scheduling logic without the need for complex interacting combinations of sub-models and custom rules, This makes our approach significantly faster and simpler to operate than existing approaches to synthesise or anonymise schedule data. We additionally contribute a novel schedule representation and a comprehensive evaluation framework. We evaluate a range of schedule encoding and deep model architecture combinations. The evaluation shows our approach can rapidly generate large, diverse, novel, and realistic synthetic samples of activity schedules.
@@
We contribute a novel approach to synthesising human activity schedules using deep generative modelling. Specifically, we demonstrate the rapid synthesis of large, realistic samples of human activity schedules using Variational Auto-Encoders.
We propose a novel activity schedule encoding, using a continuous encoding of activity durations.
We present a comprehensive framework for evaluating the quality of model outputs. We use human interpretable distance metrics such that users can reasonably assess the suitability of models for various tasks.
@@ 1 Introduction
Our definition of activity schedules can be decomposed into: (i) participations - if or how many times an individual chooses to undertake certain activities, and (ii) timings - when and for how long to undertake the chosen activities. Note, we do not consider other aspects such as activity locations or associated trips and their attributes, such as mode choices.
The participation and timing of multiple activities for an individual create a high-dimensional object with complex joint distributions. The complex distributions of activity schedules are the result of the real-world processes by which people plan and undertake their days. This process includes the consideration of (i) physical and temporal constraints, such as not being able to partake in multiple activities at the same time or fitting all activities into a finite time budget, (ii) personal and social constraints, such as arriving at work on time or needing to buy food, (iii) individual preferences, such as the desire to participate in leisure activities or preferences towards different activity start-times, and (iv) interactions with others, such as coordinating shared activities in a household or via congestion.
@@
We demonstrate a novel approach for activity schedule synthesis using deep generative machine learning (ML). Our approach allows for the simultaneous synthesis of schedules, combining choices that would typically be made separately. Additionally, deep generative ML explicitly models the variation in observed data, without the need for conditional information. This results in three key benefits: (i) simultaneous synthesis potentially allows more realistic interaction of choice components, (ii) deep generative ML maintains diversity in output sequences, and (iii) in combination, the approach is simple and fast to apply. We comprehensively evaluate the quality of output activity schedules both individually and in aggregate.
A major current limitation of our approach is that we do not consider conditionality, such as in a typical behavioural modelling framework, for example, we cannot model responses to changes in socio-economic attributes or accessibility. This limits the application to where conditionality is not required, such as: (i) synthesis, (ii) data anonymisation, and (iii) diverse up-sampling.
Ultimately, this work is intended to demonstrate and progress generative modelling approaches in this domain, to enable their introduction into models and frameworks with conditionality in future. We share experimentation with data encodings and model architectures to assist future approaches using deep ML and generative approaches. Code for this research is available from Caveat [^1] https://github.com/big-ucl/caveat.
@@ 1.1 Activity Scheduling
1.2 Machine Learning Approaches to Activity Scheduling
@@
Nayak and Pandit (2023) consider activity sequences (participations) across multiple days as a classification problem; comparing the performance of various bagging and boosting models for this task. They simplify the activity scheduling problem by identifying a reduced set of common day tours and then treating scheduling as a discriminative supervised learning problem. Hafezi et al. (2021) use Random Forests to predict activity start times and durations given person attributes. They then use heuristic rules to assemble these activity times and durations into schedules. Time is discretised into various bin sizes from 10 to 360 minutes. In such discriminative ML approaches, evaluation is made by reporting classification accuracy for various predefined clusters against ground-truth data. The use of accuracy for evaluation makes the impact of the approaches on a broader scheduling framework difficult to assess.
@@ 1.3 Activity Schedule Evaluation
A common evaluation approach involves analysing aggregate activity participation rates by time bin. Figure 1 shows this distribution for 2022 UK National Travel Survey data. Such evaluation considers total or relative participation in activities by time of day. However, by aggregating individuals’ choices, such evaluation can hide potentially unrealistic disaggregate behaviours. More generally, activity schedules are high-dimensional objects. Therefore more detailed analysis is needed to validate the realism of individual behaviours.
Figure 1: UK National Travel Survey (2021) Agg. Activity Participation by Time Bin
@@ 1.4 Deep Generative Modelling
The above approaches use or estimate likelihood as a training objective. Generative Adversarial Networks (GANs), originally proposed by Goodfellow et al. (2014), discard this requirement, instead using a sample-based objective using a generator and discriminator architecture. This approach is extremely expressive, allowing the generation of highly realistic data. However, GANs do not estimate density and, as such, the likelihood of generated samples is not modelled. Training GANs is also difficult, as the generator-discriminator architecture can be unstable and prone to mode collapse.
@@ 1.5 Evaluating Deep Generative Models
DGMs are often found to exhibit a trade-off between individual sample quality and overall density estimation, described as precision and recall by Sajjadi et al. (2018) and as fidelity and diversity by Naeem et al. (2020). In some use cases, such as image generation, sample quality is critical, and evaluation of density estimation is limited to ensuring there is sufficient diversity in generated outputs.
@@ 1.6 Generative ML in Transport Modelling
Deep generative models have also been used to generate activity schedules. Liao et al. (2024) discretise schedules into 15-minute steps. They use a complex deep autoregressive model to generate schedules and associated activity locations conditional on a range of socio-economic attributes for members of households. They evaluate generated schedules by comparing a small set of marginal distributions and demonstrate the generated schedules in a downstream traffic simulation task. An attention-based architecture outperforms various RNN architectures, although an LSTM architecture compares similarly.
Badu-Marfo et al. (2020) use a GAN to generate both synthetic attributes and sequences. They additionally train their model to allow differentiable privacy. They show promising results for marginal attribute distributions, but find that the model performs poorly at scheduling. They note that the model struggles with the complexity of learning higher-order correlations and long-range temporal dependencies.
@@
We add to the above by applying VAEs to simultaneous activity participations and scheduling. We consider novel combinations of schedule encodings and model architectures. We also introduce a comprehensive evaluation framework, designed to evaluate the suitability of generated schedules for downstream tasks.
@@ 1.7 Generative ML for Sequence Data
Bowman et al. (2015) demonstrate a sequence-generating VAE that can generate text by sampling from the latent space. They note this is effectively a combination of an auto-regressive and latent model. They show that the latent representation can learn useful information about the overall distribution of text. Variations of this architecture are used by Gregor et al. (2015) for generating images, and Roberts et al. (2018) for generating music. In some cases, models are evaluated on the Kullback–Leibler Divergence (KLD) of their latent layer and regeneration of test sequences; however, this is primarily used for model comparison. Ultimately, researchers employ qualitative evaluation of models based on intended application.
Activity schedules are sequences of joint activities and durations, which create a novel mixed data type. Schedules combine both categorical information for activities, as per text data, and continuous information for durations, as per image pixel values. Schedules also typically have a fixed total duration.
@@ 1.8 Research Contributions
  • We contribute a novel approach to synthesising human activity schedules using deep generative modelling that allows for the rapid synthesis of large, realistic samples of human activity schedules.
  • We propose a novel activity schedule encoding, using a continuous encoding of activity durations.
@@ 2.2 Activity Schedule Definition
where $n$ indexes the position in the schedule. The number of activities in the schedule (i.e. the sequence length $N$) may vary, but the total duration should equal the time period $T$, such that:
$\sum_{n}^{N}d_{n}=T.$ (2)
We use a time period $T$ of 24 hours, starting and ending at midnight. However, the approach can be trivially extended to different periods and durations.
@@ 2.3 Problem Definition
We consider a sample of schedules to be a random variable $X\subset\mathcal{X}$, where $\mathcal{X}$ is the space of all possible activity schedules. These schedules are drawn from a target distribution $P_{X}$, such that we consider them to have the same distribution, which we refer to as the real distribution. This distribution captures the process of human activity scheduling, encompassing both individual preferences and environmental constraints such as accessibility and congestion.
Let the real sample be a finite set of $N_{real}$ observed schedules:
$\mathcal{S}_{\text{real}}=\{x_{1},x_{2},\dots,x_{N_{real}}\}\subset\mathcal{X},\quad\text{with }x\sim P_{X}.$ (3)
$\mathcal{S}_{\text{real}}$ is used to train a generative model $P_{\theta}$, parametrised by $\theta$, which aims to approximate the real distribution $P_{X}$ with a learned distribution $\hat{P}_{X}$.
@@
$\mathcal{S}_{\text{syn}}=\{\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{N_{syn}}\},\quad\text{with }\hat{x}\sim\hat{P}_{X}.$ (4)
@@
Note that the distribution $\hat{P}_{X}$ includes the density, or probability, of all possible activity schedules and the joint distributions within each schedule. As per the real sample, these are for realised schedules, accounting for preferences and constraints. For all the following experiments, we choose the size of the synthetic sample ($N_{real}$) to match the size of the real sample ($N_{syn}$).
@@ 2.4 Generative Modelling with Variational Auto-encoders
Figure 2: VAE Models Template
@@
For the former, reconstruction loss (ReconLoss), measures the difference between each output schedule and the target input schedule. Reconstruction losses are described in Section 4.3. For the latter, Kullback-Leibler Divergence (KLD) loss measures the difference between the latent layer distribution and the target normal distribution. The final loss function is composed of ReconLoss and KLD, where KLD is factored using the hyperparameter $\beta$ (reported separately for each model in Table 4).
$\text{Loss}=\text{ReconLoss}+\beta\text{KLD}$
@@ 2.5 Observed Data
To artificially create structural zeros within the observed data, which we later use for quantitative sample quality evaluation, the schedules are: (i) filtered to exclude non-home-based schedules (schedules that do not start and end at home), and (ii) modified, such that consecutive home, work or education activities are combined into single activities of the same type with combined duration. Figure  3 shows example activity sequences from the UK NTS.
Figure 3: Example NTS Schedules
@@ 2.6 Experiment Design
Our evaluation framework is described in Section 5. This evaluation is used for high-level comparison of schedule encodings and model architectures in Section 6, and more detailed consideration of the best-performing models in Section 7. We supplement these results with different scenarios, considering data sample sizes in Section 8 and alternative COVID-19 behaviours in Section 9.
Table 2: Experiments Overview
@@ 4 Model Design
KLD Kullback-Leibler Divergence
@@ 4.1 Latent Block
Figure 6 presents the latent block structure in all models. Within the block, the latent layer, a vector of length six, is sampled from a normal distribution with mean $\mu$ and standard deviation $\sigma$. This is commonly referred to as the reparametrisation trick. $\mu$ and $\sigma$ are both vectors with the same size as the latent layer. $\mu$ and $\sigma$ are resized from the encoder output using fully connected linear layers. After sampling, the latent layer is resized as required using a fully connected linear layer for passing into the decoder block.
Figure 6: Latent Block

### v3

@@ Abstract
Using a deep generative machine learning approach, we synthesise human activity participations and scheduling; i.e. the choices of what activities to participate in and when. Activity schedules are a core component of many applied transport, energy, and epidemiology models. Our data-driven approach directly learns the distributions resulting from human preferences and scheduling logic without the need for complex interacting combinations of sub-models and custom rules. This makes our approach significantly faster and simpler to operate than existing approaches to synthesise or anonymise schedule data. We additionally contribute a novel schedule representation and a comprehensive evaluation framework. We evaluate a range of schedule encoding and deep model architecture combinations. The evaluation shows our approach can rapidly generate large, diverse, novel, and realistic synthetic samples of activity schedules.
@@
We contribute a novel approach to synthesising human activity schedules using deep generative modelling. Specifically, we demonstrate the rapid synthesis of large, realistic, and representative samples of human activity schedules using Variational Auto-Encoders.
We propose an activity schedule encoding, using a continuous encoding of activity durations.
We present a comprehensive framework for evaluating the quality of model outputs, focusing on distribution-level factors which are crucial for simulations. We use human interpretable distance metrics such that users can reasonably assess the suitability of models for various tasks.
@@ 1 Introduction
Our definition of activity schedules can be decomposed into: (i) participations - if or how many times an individual chooses to undertake certain activities, and (ii) timings - when and for how long to undertake the chosen activities. Note, we do not consider here other aspects such as activity locations or associated trips and their attributes, such as mode choices.
The participation and timing of multiple activities for an individual creates a high-dimensional object with complex joint distributions. The complex distributions of activity schedules are the result of the real-world processes by which people plan and undertake their days. This process includes the consideration of (i) physical and temporal constraints, such as not being able to partake in multiple activities at the same time or fitting all activities into a finite time budget, (ii) personal and social constraints, such as arriving at work on time or needing to buy food, (iii) individual preferences, such as the desire to participate in leisure activities or preferences towards different activity start-times, and (iv) interactions with others, such as coordinating shared activities in a household or via congestion.
@@
We demonstrate a novel approach for activity schedule synthesis using deep generative machine learning (ML). Our approach allows for the simultaneous synthesis of schedules, combining choices that would typically be made separately. Additionally, deep generative ML explicitly models the variation in observed data, without the need for conditional information. This results in three key benefits: (i) simultaneous synthesis potentially allows more realistic interaction of choice components, (ii) deep generative ML better models the distribution of output sequences, and (iii) in combination, the approach is simple and fast to apply. We comprehensively evaluate the quality of output activity schedules both individually and in aggregate.
We show that by incorporating a generative approach, we can better model the true distribution, or variance, of real schedules compared to existing conditional supervised learning approaches. However, a major limitation is that we cannot model conditionality, such as in a typical behavioural modelling framework. For example, we cannot model responses to changes in socio-economic attributes or accessibility. This limits the application to where conditionality is not required, such as: (i) synthesis, (ii) data anonymisation, and (iii) diverse up-sampling.
Ultimately, this work is intended to demonstrate and establish generative modelling approaches in this domain, leaving their introduction into models and frameworks with conditionality to future work. We share experimentation with data encodings and model architectures to assist future approaches using deep ML and generative approaches. Code for this research is available from the Caveat GitHub repository [^1] https://github.com/big-ucl/caveat.
@@ 1.1 Activity Scheduling
1.2 Supervised Discriminative Machine Learning Approaches to Activity Scheduling
@@
Nayak and Pandit (2023) consider activity sequences (participations) across multiple days as a classification problem; comparing the performance of various bagging and boosting models for this task. They simplify the activity scheduling problem by identifying a reduced set of common day tours and then treating scheduling as a supervised discriminative (or conditional) learning problem. Hafezi et al. (2021) use Random Forests to predict activity start times and durations given person attributes. They then use heuristic rules to assemble these activity times and durations into schedules. Time is discretised into various bin sizes from 10 to 360 minutes. In such discriminative ML approaches, evaluation is made by reporting classification accuracy for various predefined clusters against ground-truth data. The use of accuracy for evaluation makes it difficult to assess the impact of the approaches on a broader scheduling framework.
@@ 1.3 Activity Schedule Evaluation
A common evaluation approach involves analysing aggregate activity participation rates by time bin. Figure 1 shows this distribution for 2023 UK National Travel Survey data. Such evaluation considers total or relative participation in activities by time of day. However, by aggregating individuals’ choices, such evaluation can hide potentially unrealistic disaggregate behaviours. More generally, activity schedules are high-dimensional objects. Therefore more detailed analysis is needed to validate the realism of individual behaviours.
Figure 1: UK National Travel Survey (2023) Activity Frequencies: Aggregate activity participation by 10-minute time bin
@@ 1.4 Deep Generative Modelling
The above approaches use or estimate likelihood as a training objective. Generative Adversarial Networks (GANs), originally proposed by Goodfellow et al. (2014), discard this requirement, instead using a sample-based objective using a generator and discriminator architecture. This approach is extremely expressive, allowing the generation of highly realistic data. However, GANs do not estimate density and, as such, the likelihood of generated samples is not modelled. Training GANs is also difficult, as the generator-discriminator architecture can be unstable and prone to mode collapse, where the generator produces a limited or repetitive set of outputs.
@@ 1.5 Evaluating Deep Generative Models
DGMs are often found to exhibit a trade-off between individual sample quality and aggregate density estimation, described as precision and recall by Sajjadi et al. (2018) and as fidelity and diversity by Naeem et al. (2020). In some use cases, such as image generation, sample quality is critical, and evaluation of density estimation is limited to ensuring there is sufficient diversity in generated outputs.
@@ 1.6 Generative ML in Transport Modelling
Deep generative models have also been used to model activity schedules. Badu-Marfo et al. (2020) use a GAN to generate both synthetic attributes and sequences. They show promising results for marginal attribute distributions, but find that the model performs poorly at scheduling. They note that the model struggles with the complexity of learning higher-order correlations and long-range temporal dependencies.
Liao et al. (2024) discretise schedules into 15-minute steps. They use a complex deep autoregressive model to generate schedules and associated activity locations conditional on a range of socio-economic attributes for members of households. They evaluate generated schedules by comparing a small set of marginal distributions and demonstrate the generated schedules in a downstream traffic simulation task. An attention-based architecture outperforms various RNN architectures, although an LSTM architecture compares similarly.
@@
1.7 Summary of Discriminative and Generative ML Approaches for Activity Scheduling
Table 2 presents an overview of existing ML approaches to activity scheduling, combining both the supervised discriminative learning approaches introduced in Section 1.2 and the generative learning approaches in Section 1.6. All but one of the existing approaches are discriminative or conditional-generative, modelling part or all the variance in observed schedules using explanatory variables such as sociodemographics. They are therefore trained using pairs of attributes and schedules, or equivalent, using a supervised learning approach. Only Badu-Marfo et al. (2020) use a purely self-supervised generative approach.
Table 2: Summary of Existing Machine Learning Activity Scheduling Approaches
Model Name Output Approach
Nayak and Pandit (2023) activity participations supervised discriminative
Hafezi et al. (2021) activity timings supervised discriminative
Koushik et al. (2023) activity schedules supervised discriminative
Liao et al. (2024) activity schedules & locations supervised conditional-generative
Kim et al. (2022) activity types supervised conditional-generative
Badu-Marfo et al. (2020) activity schedules self-supervised generative
We provide a purely generative approach to activity scheduling that does not require explanatory data to model variance in scheduling. We add to the generative approach by Badu-Marfo et al. (2020) by considering a VAE approach with the goal of better modelling the true distribution of real schedules, as well as novel combinations of schedule encodings and model architectures. We also introduce a comprehensive evaluation framework, designed to evaluate the suitability of generated schedules for downstream tasks.
@@ 1.8 Generative ML for Sequence Data
Bowman et al. (2015) demonstrate a sequence-generating VAE that can generate text by sampling from the latent space. They note this is effectively a combination of an auto-regressive and latent model. They show that the latent representation can learn useful information about the overall distribution of text. Variations of this architecture are used by Gregor et al. (2015) for generating images, and Roberts et al. (2018) for generating music. In some cases, models are evaluated on the Kullback–Leibler Divergence ($D_{KL}$) of their latent layer and regeneration of test sequences; however, this is primarily used for model comparison. Ultimately, researchers employ qualitative evaluation of models based on intended application.
Activity schedules are sequences of activity types and durations, combining categorical information, as per text data, and continuous information, as per image pixel values. Schedules also typically have a fixed total duration. We build on this observation by proposing and testing a mixed-type schedule representation.
@@ 1.9 Research Contributions
  • We demonstrate an approach to synthesising human activity schedules using VAEs that can rapidly generate realistic samples of human activity schedules.
  • We propose an activity schedule encoding, using a continuous encoding of activity durations.
@@ 2.2 Activity Schedule Definition
where $n$ indexes the position in the schedule. Activity types are limited to a finite set $\mathcal{A}$. The number of activities in the schedule (i.e. the sequence length $N$) may vary, but the total duration should equal the time period $T$, such that:
$\sum_{n=1}^{N}d_{n}=T.$ (2)
We use a time period $T$ of 24 hours, starting and ending at midnight. However, the approach can be extended to different periods and durations.
@@ 2.3 Problem Definition
We consider schedules to be a random variable $X\subset\mathcal{X}$, where $\mathcal{X}$ is the space of all possible activity schedules. These schedules are drawn from a target distribution $P_{X}$, such that we consider them to have the same distribution, which we refer to as the real distribution. This distribution captures the process of human activity scheduling, encompassing both individual preferences and environmental constraints such as accessibility and congestion.
Let the real sample be composed of $N_{real}$ observed schedules:
$X_{\text{real}}=(x_{1},x_{2},\dots,x_{N_{real}})\subset\mathcal{X},\quad\text{with }x\sim P_{X}.$ (3)
$X_{\text{real}}$ is used to train a generative model $P_{\theta}$, parametrised by $\theta$, which aims to approximate the real distribution $P_{X}$ with a learned distribution $\hat{P}_{X}$.
@@
$X_{\text{syn}}=(\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{N_{syn}}),\quad\text{with }\hat{x}\sim\hat{P}_{X}.$ (4)
@@
Note that the distribution $\hat{P}_{X}$ includes the density, or probability, of all possible activity schedules and the joint distributions within each schedule. As per the real sample, these are for realised schedules, accounting for preferences and constraints. For all the following experiments, we choose the size of the synthetic sample ($N_{syn}$) to match the size of the real sample ($N_{real}$).
@@ 2.4 Generative Modelling with Variational Auto-encoders
Figure 2: VAE Model Template
@@
VAE encoder architectures use a stochastic process, commonly called the reparametrisation trick, to model the random distribution $Z$. This is described in more detail in Section 4.1. Although not specifically required for the VAE approach, we use deterministic decoder architectures throughout, such that, for generation, schedules depend on sampling $Z$ only.
@@
For the former, reconstruction loss (ReconLoss), measures the difference between each output schedule and the target input schedule. Reconstruction losses are described in Section 4.3. For the latter, Kullback-Leibler Divergence ($D_{KL}$) loss measures the difference between the latent layer distribution and the target normal distribution. The final loss function is composed of ReconLoss and $D_{KL}$, where $D_{KL}$ is factored using the hyperparameter $\beta$ (reported separately for each model in Table 6).
$\text{Loss}=\text{ReconLoss}+\beta D_{KL}$
@@ 2.5 Observed Data
We modify the observed data to impose structural constraints, which we later use for quantitative sample quality evaluation. The schedules are: (i) filteredto exclude non-home-based schedules (schedules that do not start and end at home), and (ii) modified, such that consecutive home, work or education activities are combined into single activities of the same type with combined duration .
These constraints can be considered as structural zeros, for example, we might expect all schedules to start and end at home and not include trips from home to home. If so, imposing them can be considered a data cleaning exercise. However, such schedules are observed, although rarely. Therefore, more pragmatically and independent of any specific broader framework requirements, these constraints can be considered as ways to evaluate model compliance - whether the models can be constrained to structure schedules in specific ways or not. Figure  3 shows example activity sequences from the UK NTS.
Figure 3: Example Schedules from the UK National Travel Survey (2023)
@@ 2.6 Experiment Design
Our evaluation framework is described in Section 5. This evaluation is used for high-level comparison of schedule encodings and model architectures in Section 6, and more detailed consideration of the best-performing models in Section 7. We supplement these results with different scenarios, considering data sample sizes in Section 8.1 and alternative scenarios in Section 8.2.
Table 3: Summary of VAE Model Architectures
@@ 3.1 Discrete Schedule Encoding
The 10-minute time step is primarily chosen to be consistent with prior work, such as by Koushik et al. (2023). This is additionally a pragmatic choice to allow for precise temporal predictions and the representation of short activities that might otherwise be omitted for longer step sizes. Table 4 shows ranked evaluations (evaluation metrics are detailed later in Section 5) for the DiscCNN model using various step sizes between 5 minutes and 2 hours. This experimentation suggests that the 10-minute step size provides a good trade-off between the evaluation metrics.
Table 4: Discrete Step Size Experiment - Evaluation Rankings
5min 10min 15min 20min 30min 60min 120min
Density Estimation
Participations 6 2 4 3 5 1 7
Transitions 7 2 3 4 5 1 6
Timing 6 1 2 4 5 3 7
Validity 1 1 1 1 1 1 1
Creativity 2 1 3 4 5 6 7
@@ 4 Model Design
$D_{KL}$ Kullback-Leibler Divergence
@@ 4.1 Latent Block
The VAE encoders model the latent vector $Z$ as a random variable. This cannot be directly modelled as a random sample within a VAE, because such an operation would be non-differentiable. Instead, VAEs model the latent space using learnt means and variances, allowing for differentiation; this is commonly referred to as the reparametrisation trick and is described in more detail by Kingma and Welling (2013).
All models use the same latent block structure shown in Figure 6. Within the block, the latent layer, a vector of length six, is sampled from a normal distribution with mean $\mu$ and standard deviation $\sigma$. $\mu$ and $\sigma$ are both vectors with the same size as the latent layer. $\mu$ and $\sigma$ are resized from the encoder output using fully connected linear layers. After sampling, the latent layer is resized as required using a fully connected linear layer for passing into the decoder block.
Figure 6: Latent Block Architecture